### v1

@@ 3.1 Simulation settings
To fully characterize marine conditions where the response variables of regression models often exhibit distinct spatial structures in geoscientific research, we consider two-dimensional spatial data in the simulations, which is commonly considered in marine across longitude and latitude coordinates or longitude and depth. Without losing generality, we assume a $[0,1]\times[0,1]$ spatial domain with $n=1,000$ randomly distributed observation points $s_{1},\dots,s_{n}\in\mathbb{R}^{2}$.
@@
In the generating process, parameter $\phi$ controls the spatial heterogeneity. Then the spatial features $x_{1}(s_{i})$ and $x_{2}(s_{i})$ are generated from $z_{1}(s_{i})$ and $z_{2}(s_{i})$ by linear transformation. Specifically, we set $x_{1}(s_{i})=z_{1}(s_{i})$ and $x_{2}(s_{i})=rz_{1}(s_{i})+\sqrt{1-r^{2}}z_{2}(s_{i})$, where the parameter $r$ controls the level of collinearity. We set $r=0.75$ for a strong correlation design and generate five datasets with $\phi=0.2,0.4,0.6,0.8,1$ for different levels of spatial correlation, which are labeled as “DGP1”, $\cdots$, “DGP5”, respectively. The true values of the parameters $\beta_{1},\beta_{2},\beta_{3}$ are set as shown in Figure 1(a)-(c) respectively, where different clustering patterns are adopted for three different parameters.
To quantify the performance of estimations for dimension $k\in[p]$ and location $i\in[n]$, we focus on the three measures with $T=100$ independent replications: (1) Root Mean Squared Error (RMSE), which is defined as $\text{RMSE}_{k}^{(i)}=\left[\sum\limits_{t=1}^{T}\left(\hat{\beta}_{k}^{t}(s_{i})-\beta_{k}(s_{i})\right)^{2}/T\right]^{1/2}$ (2) Mean Absolute Error (MAE), which is defined as $\text{MAE}_{k}^{(i)}=\sum\limits_{t=1}^{T}|\hat{\beta}_{k}^{t}(s_{i})-\beta_{k}(s_{i})|/T$; (3) Standard Deviation (SD), which is defined as $\text{SD}_{k}^{(i)}=\left[\sum\limits_{t=1}^{T}\left(\hat{\beta}_{k}^{t}(s_{i})-\bar{\hat{\beta}}_{k}(s_{i})\right)^{2}/T\right]^{1/2}$, where $\bar{\hat{\beta}}_{k}(s_{i})=\sum\limits_{t=1}^{T}\hat{\beta}_{k}^{t}(s_{i})/T$. Smaller values in RMSE, MAE and SD indicate more accurate and stable estimations for parameters in model (12). As we have in total $np$ parameters to estimate, the mean values of the aforementioned measures across all spatial points $i\in[n]$ for each dimension $k\in[p]$ are computed and reported in Table 1.
To assess the performance in detecting discrete spatial heterogeneity in each dimension $k\in[p]$, we consider four widely used measures for clustering: Rand Index (RI), Adjusted Rand Index (ARI), Silhouette Coefficient (SC) and Calinski-Harabasz Index (CHI). Higher values of RI, ARI, SC and CHI reflect a greater consistency between the detected clusters and the true group labels. Detailed definitions and further discussion of these metrics are included in A.
@@ 3.2 Simulation results
Table 1: The comparison of the performance in estimation with 5 data-generating processes
Evaluation Parameter Method DGP1 DGP2 DGP3 DGP4 DGP5
RMSE $\beta_{1}$ GWR 3.0521 4.2638 5.1742 5.9329 6.7977
SCC 0.8612 1.0648 1.2451 1.3882 1.5690
FLAT 0.8090 0.9800 1.1503 1.2817 1.4446
$\beta_{2}$ GWR 3.0344 4.2502 5.1628 5.9215 6.8203
SCC 0.9390 1.1372 1.3132 1.4588 1.6872
FLAT 0.8894 1.0454 1.2160 1.3582 1.5665
$\beta_{3}$ GWR 3.0311 4.5394 5.6082 6.4590 7.5956
SCC 0.9015 1.1835 1.4219 1.5988 1.8425
FLAT 0.8502 1.0845 1.3177 1.4848 1.7280
MAE $\beta_{1}$ GWR 2.3340 3.2390 3.9096 4.4751 5.0945
SCC 0.5856 0.7330 0.8726 0.9953 1.1371
FLAT 0.5370 0.6645 0.7906 0.8984 1.0293
$\beta_{2}$ GWR 2.3247 3.2121 3.8829 4.4486 5.0979
SCC 0.6520 0.8019 0.9443 1.0641 1.2616
FLAT 0.5940 0.7191 0.8520 0.9684 1.1459
$\beta_{3}$ GWR 2.0214 2.8968 3.5364 4.0649 4.7283
SCC 0.5698 0.7448 0.8997 1.0234 1.2073
FLAT 0.5395 0.6884 0.8280 0.9445 1.1196
SD $\beta_{1}$ GWR 2.9521 4.1770 5.0925 5.8530 6.7234
SCC 0.7213 1.4430 1.0591 1.1813 1.3341
FLAT 0.6888 0.8401 0.9865 1.0991 1.2338
$\beta_{2}$ GWR 2.9500 4.1666 5.0833 5.8433 6.7586
SCC 0.7721 1.7088 1.0781 1.1950 1.3522
FLAT 0.7501 0.8764 1.0102 1.1270 1.2729
$\beta_{3}$ GWR 2.7790 4.3506 5.4442 6.3084 7.4337
SCC 0.7644 1.8384 1.2170 1.3703 1.5622
FLAT 0.7151 0.9165 1.1237 1.2681 1.4605
@@
The comparison based on the measures of clustering is shown in Table 2. As shown in the first row of Figure 1, true spatial parameters have spatial clustering patterns. We conduct the identical standard DBSCAN clustering algorithm with the estimations by GWR, SCC and FLAT and calculate the RI, ARI, SC, and CHI, which measure the similarity between the discrete spatial heterogeneity and the detected clusters of coefficients in different dimensions. Thus, more accurate clustering results will indicate better capture of discrete spatial heterogeneity. As the Table 2, FLAT achieves better performance with higher RI, ARI, SC and CHI compared to the classic methods GWR and SCC, which indicates the effectiveness of FLAT in detecting the discrete spatial heterogeneity in practice. Similar conclusions are found with other data-generating processes, and we include the additional numerical results in B.
Table 2: Evaluations based on clustering results of estimated parameters $\hat{\bm{\beta}}$ by different methods
Data-generating process Parameter Method RI ARI SC CHI(×$\times 10^{4}$)
DGP1 $\beta_{1}$ GWR 0.6411 0.2453 0.4537 0.2041
SCC 0.9117 0.7898 0.8685 2.3816
FLAT 0.9180 0.8051 0.8791 3.3122
$\beta_{2}$ GWR 0.7446 0.2769 0.4163 0.5611
SCC 0.9185 0.7490 0.8378 3.1291
FLAT 0.9289 0.7804 0.8577 4.0711
$\beta_{3}$ GWR 0.8084 0.3576 0.5163 0.9952
SCC 0.9346 0.7709 0.8713 4.3364
FLAT 0.9370 0.7818 0.8835 4.7318
DGP5 $\beta_{1}$ GWR 0.6101 0.1260 0.1879 0.3858
SCC 0.8056 0.5650 0.7494 0.7449
FLAT 0.8227 0.5951 0.7680 1.0421
$\beta_{2}$ GWR 0.6760 0.1487 0.1363 0.7457
SCC 0.8307 0.5023 0.7572 1.6862
FLAT 0.8471 0.5386 0.7754 2.3999
$\beta_{3}$ GWR 0.7176 0.2061 0.2091 1.3860
SCC 0.8640 0.5660 0.7871 2.0479
FLAT 0.8718 0.5892 0.8066 2.1975
@@ 4 Real data analysis
Ocean stratification is fundamentally tied to T-S gradients. Warmer or fresher (less dense) waters tend to overlie colder or saltier (denser) layers, exemplified by Antarctic Intermediate Water. This stratification shapes vertical mixing, nutrient fluxes, and biogeochemical cycles. To quantify spatial gradients, we adopt a numerical method called the spatial difference quotient, which measures the rate of change across discrete spatial datasets-analogous to the derivative in continuous calculus. We introduce these concepts and corresponding formulas in C.
@@ 4.2 Estimation results
7
13
RMSE SC CHI(×$\times 10^{3}$) RMSE SC CHI(×$\times 10^{3}$)
3
5
7
9
11
13
@@ 4.3 Analysis of the T-S relationship at sea surface layers
We further calculated the Spatial Difference Quotient (SDQ) of the sea surface T-S relationship based on the estimated coefficients from Figure 5, as shown in Figure 5. The SDQ can be interpreted as a numerical gradient of the T-S relationship at each spatial location; more details on its definition are provided in C. As shown in Figure 5, FLAT effectively captures the spatial variation patterns of T-S relationships. Regions with larger difference quotients are primarily located near landmasses and the equator, particularly near two estuarine zones-the eastern Atlantic coast of North America and northeastern South America. In the equatorial zone, especially at 5°N latitude and between 20°E and 50°E longitude, the T-S coefficient approaches zero, while the SDQ remains relatively high. This pattern likely arises from the distinct marine conditions in equatorial waters: higher temperatures result in reduced density, while intense solar radiation drives strong evaporation and elevated salinity. The counteracting effects of temperature and salinity yield a near-zero T-S coefficient, and the sharp local contrasts contribute to the observed high spatial gradients, reflecting a dynamic T-S compensation regime.
@@ 5 Conclusion and Discussion
Appendix A The evaluations of clustering

### v2

@@ 3.1 Simulation settings
To fully characterize marine conditions where the response variables of regression models often exhibit distinct spatial structures in geoscientific research, we consider two-dimensional spatial data in the simulations, which is commonly considered in marine across longitude and latitude coordinates or longitude and depth. Without losing generality, we assume a $0,1]\times[0,1]$ spatial domain with $n=1,000$ randomly distributed observation points $s_{1},\dots,s_{n}\in\mathbb{R}^{2}$.
@@
In the generating process, parameter $\phi$ controls the spatial heterogeneity. Then the spatial features $x_{1}(s_{i})$ and $x_{2}(s_{i})$ are generated from $z_{1}(s_{i})$ and $z_{2}(s_{i})$ by linear transformation. Specifically, we set $x_{1}(s_{i})=z_{1}(s_{i})$ and $x_{2}(s_{i})=rz_{1}(s_{i})+\sqrt{1-r^{2}}z_{2}(s_{i})$, where the parameter $r$ controls the level of collinearity. We set $r=0.75$ for a strong correlation design and generate five datasets with $\phi=0.2,0.4,0.6,0.8,1$ for different levels of spatial correlation. The true values of the parameters $\beta_{1},\beta_{2},\beta_{3}$ are set as shown in Figure 1(a)-(c) respectively, where different clustering patterns are adopted for three different parameters.
To quantify the performance of estimations for dimension $k\in[p]$ and location $i\in[n]$, we focus on the three measures with $T=100$ independent replications: (1) Root Mean Squared Error (RMSE), which is defined as $\text{RMSE}_{k}^{(i)}=[\sum_{t=1}^{T}\{\hat{\beta}_{k}^{t}(s_{i})-\beta_{k}(s_{i})\}^{2}/T]^{1/2}$ (2) Mean Absolute Error (MAE), which is defined as $\text{MAE}_{k}^{(i)}=\sum_{t=1}^{T}|\hat{\beta}_{k}^{t}(s_{i})-\beta_{k}(s_{i})|/T$; (3) Standard Deviation (SD), which is defined as $\text{SD}_{k}^{(i)}=[\sum_{t=1}^{T}\{\hat{\beta}_{k}^{t}(s_{i})-\bar{\hat{\beta}}_{k}(s_{i})\}^{2}/T]^{1/2}$, where $\bar{\hat{\beta}}_{k}(s_{i})=\sum_{t=1}^{T}\hat{\beta}_{k}^{t}(s_{i})/T$. Smaller values in RMSE, MAE and SD indicate more accurate and stable estimations for parameters in model (6). As we have in total $np$ parameters to estimate, the mean values of the aforementioned measures across all spatial points $i\in[n]$ for each dimension $k\in[p]$ are computed and reported in Table 1.
To assess the performance in detecting discrete spatial heterogeneity in each dimension $k\in[p]$, we consider four widely used measures for clustering: Rand Index (RI), Adjusted Rand Index (ARI), Silhouette Coefficient (SC) and Calinski-Harabasz Index (CHI). Higher values of RI, ARI, SC and CHI reflect a greater consistency between the detected clusters and the true group labels. Detailed definitions and further discussion of these metrics are included in 5.2.
@@ 3.2 Simulation results
Table 1: The comparison of estimation performance with different levels of spatial correlation: The mean of the different measures across all spatial points $i\in[n]$ are computed and reported for each individual parameter with different level of spatical correlation.
GWR SCC FLAT
Spatial Correlation $\phi$ Evaluation $\beta_{1}$ $\beta_{2}$ $\beta_{3}$ $\beta_{1}$ $\beta_{2}$ $\beta_{3}$ $\beta_{1}$ $\beta_{2}$ $\beta_{3}$
0.2 RMSE 3.0521 3.0344 3.0311 0.8612 0.9390 0.9015 0.8090 0.8894 0.8502
MAE 2.3340 2.3247 2.0214 0.5856 0.6520 0.5698 0.5370 0.5940 0.5395
SD 2.9521 2.9500 2.7790 0.7213 0.7721 0.7644 0.6888 0.7501 0.7151
0.4 RMSE 4.2638 4.2502 4.5394 1.0648 1.1372 1.1835 0.9800 1.0454 1.0845
MAE 3.2390 3.2121 2.8968 0.7330 0.8019 0.7448 0.6645 0.7191 0.6884
SD 4.1770 4.1666 4.3506 1.4430 1.7088 1.8384 0.8401 0.8764 0.9165
0.6 RMSE 5.1742 5.1628 5.6082 1.2451 1.3132 1.4219 1.1503 1.2160 1.3177
MAE 3.9096 3.8829 3.5364 0.8726 0.9443 0.8997 0.7906 0.8520 0.8280
SD 5.0925 5.0833 5.4442 1.0591 1.0781 1.2170 0.9865 1.0102 1.1237
0.8 RMSE 5.9329 5.9215 6.4590 1.3882 1.4588 1.5988 1.2817 1.3582 1.4848
MAE 4.4751 4.4486 4.0649 0.9953 1.0641 1.0234 0.8984 0.9684 0.9445
SD 5.8530 5.8433 6.3084 1.1813 1.1950 1.3703 1.0991 1.1270 1.2681
1 RMSE 6.7977 6.8203 7.5956 1.5690 1.6872 1.8425 1.4446 1.5665 1.7280
MAE 5.0945 5.0979 4.7283 1.1371 1.2616 1.2073 1.0293 1.1459 1.1196
SD 6.7234 6.7586 7.4337 1.3341 1.3522 1.5622 1.2338 1.2729 1.4605
@@
The comparison based on the measures of clustering is shown in Table 2. As shown in the first row of Figure 1, true spatial parameters have spatial clustering patterns. We conduct the identical standard DBSCAN clustering algorithm with the estimations by GWR, SCC and FLAT and calculate the RI, ARI, SC, and CHI, which measure the similarity between the discrete spatial heterogeneity and the detected clusters of coefficients in different dimensions. Thus, more accurate clustering results will indicate better capture of discrete spatial heterogeneity. FLAT achieves better performance with higher RI, ARI, SC and CHI compared to the classic methods GWR and SCC, which indicates the effectiveness of FLAT in detecting the discrete spatial heterogeneity in practice.
Table 2: Evaluations of estimation performance based on clustering with different level of spatial correlation: RI represents the Rand index, ARI represents the adjusted Rand index, SC represents the Silhouette coefficient, and CHI represents Calinski-Harabasz index and scaled by $\times 10^{4}$.
GWR SCC FLAT
Spatial Correlation $\phi$ Evaluation $\beta_{1}$ $\beta_{2}$ $\beta_{3}$ $\beta_{1}$ $\beta_{2}$ $\beta_{3}$ $\beta_{1}$ $\beta_{2}$ $\beta_{3}$
0.2 RI 0.6411 0.7446 0.8084 0.9117 0.9185 0.9346 0.9180 0.9289 0.9370
ARI 0.2453 0.2769 0.3576 0.7898 0.7490 0.7709 0.8051 0.7804 0.7818
SC 0.4537 0.4163 0.5163 0.8685 0.8378 0.8713 0.8791 0.8577 0.8835
CHI 0.2041 0.5611 0.9952 2.3816 3.1291 4.3364 3.3122 4.0711 4.7318
0.4 RI 0.6220 0.7210 0.7800 0.8816 0.8942 0.9002 0.8932 0.9074 0.9107
ARI 0.1843 0.2173 0.2933 0.7280 0.6826 0.6753 0.7519 0.7176 0.7056
SC 0.3160 0.3331 0.4142 0.8373 0.8137 0.8432 0.8492 0.8307 0.8441
CHI 0.2232 0.5846 1.0047 1.5141 2.4471 2.9139 2.0688 3.2222 3.6620
0.6 RI 0.6298 0.7027 0.7619 0.8525 0.8732 0.8832 0.8697 0.8904 0.8959
ARI 0.1702 0.1853 0.2607 0.6607 0.6185 0.6296 0.6946 0.6672 0.6589
SC 0.2606 0.2720 0.3454 0.7910 0.7928 0.8217 0.8236 0.8204 0.8281
CHI 0.2975 0.6311 1.0801 1.2801 2.0624 2.4289 1.8807 2.6647 2.8204
0.8 RI 0.6206 0.6968 0.7467 0.8286 0.8559 0.8661 0.8516 0.8674 0.8799
ARI 0.1481 0.1681 0.2370 0.6071 0.5701 0.5871 0.6545 0.6030 0.6150
SC 0.2246 0.2144 0.2862 0.7702 0.7715 0.7931 0.7990 0.7841 0.8172
CHI 0.3315 0.7392 1.1285 1.0753 1.8524 2.0760 1.5322 2.5714 2.4064
1 RI 0.6101 0.6760 0.7176 0.8056 0.8307 0.8640 0.8227 0.8471 0.8718
ARI 0.1260 0.1487 0.2061 0.5650 0.5023 0.5660 0.5951 0.5386 0.5892
SC 0.1879 0.1363 0.2091 0.7494 0.7572 0.7871 0.7680 0.7754 0.8066
CHI 0.3858 0.7457 1.3860 0.7449 1.6862 2.0479 1.0421 2.3999 2.1975
@@ 4 Real Data Analysis
Ocean stratification is fundamentally tied to T-S gradients. Warmer or fresher (less dense) waters tend to overlie colder or saltier (denser) layers, exemplified by Antarctic Intermediate Water. This stratification shapes vertical mixing, nutrient fluxes, and biogeochemical cycles. To quantify spatial gradients, we adopt a numerical method called the spatial difference quotient, which measures the rate of change across discrete spatial datasets-analogous to the derivative in continuous calculus. We introduce these concepts and corresponding formulas in 5.3.
@@ 4.2 Estimation results
RMSE SC CHI($\times 10^{3}$) RMSE SC CHI($\times 10^{3}$)
@@ 4.3 Analysis of the T-S relationship at sea surface layers
We further calculated the Spatial Difference Quotient (SDQ) of the sea surface T-S relationship based on the estimated coefficients from Figure 5, as shown in Figure 5. The SDQ can be interpreted as a numerical gradient of the T-S relationship at each spatial location; more details on its definition are provided in 5.3. As shown in Figure 5, FLAT effectively captures the spatial variation patterns of T-S relationships. Regions with larger difference quotients are primarily located near landmasses and the equator, particularly near two estuarine zones-the eastern Atlantic coast of North America and northeastern South America. In the equatorial zone, especially at 5°N latitude and between 20°E and 50°E longitude, the T-S coefficient approaches zero, while the SDQ remains relatively high. This pattern likely arises from the distinct marine conditions in equatorial waters: higher temperatures result in reduced density, while intense solar radiation drives strong evaporation and elevated salinity. The counteracting effects of temperature and salinity yield a near-zero T-S coefficient, and the sharp local contrasts contribute to the observed high spatial gradients, reflecting a dynamic T-S compensation regime.
@@ 5 Conclusion and Discussion
Appendix
5.1 FLAT algorithm details
As introduced in Section 2, we set the adaptive weights as the spatial distance between corresponding spatial locations, then the regularization terms based on adaptive-minimum-spanning tree can be further formulated as
where $\mathbb{E}$ represents the edge set derived from the minimum spanning tree, and $\mathcal{N}_{i}$ denotes the neighboring nodes of node $i$ in the minimum spanning tree. With some derivation, the regularization for spatial heterogeneity can be rewritten as
$P_{\lambda_{1}}(\bm{\beta})=\lambda_{1}\sum_{k=1}^{p}\|\bm{\pi}\bm{H}\bm{\beta}_{k}\|_{1},$ (10)
based on the definition of matrix $\bm{H}$, where the sparse weight matrix $\bm{\pi}\in\mathbb{R}^{(n-1)\times(n-1)}$ contains all the weights of n1n-1 potential edges. For simplicity, we use matrix notation and rewrite the objective function of FLAT (5) as
$L_{\text{FLAT}}(\bm{\beta})=\frac{1}{2}\|\mathbf{y}-\bm{x}\bm{\beta}\|_{2}^{2}+\lambda_{1}\sum_{k=1}^{p}\|\bm{\pi}\bm{H}\bm{\beta}_{k}\|_{1}+\lambda_{2}\sum_{k=1}^{p}\|\bm{\beta}_{k}\|_{1},$ (11)
where $\mathbf{y}=[y(s_{1}),...,y(s_{n})]^{\top}\in\mathbb{R}^{n}$ represents the vector of response values for $n$ locations. $\bm{x}\triangleq[\text{diag}(\bm{x}_{1}),...,\text{diag}(\bm{x}_{p})]\in\mathbb{R}^{n\times np}$, where $\bm{x}_{k}$ represents the $k$-th covariate vector of length $n$, and the observation vector is $\bm{x}_{k}\triangleq(x_{k}(s_{1}),...,x_{k}(s_{n}))^{\top}\in\mathbb{R}^{n}$ is the corresponding design matrix with spatial feature. We denote $\bm{\beta}=[\bm{\beta}^{\top}(s_{1}),...,\bm{\beta}^{\top}(s_{n})]^{\top}\in\mathbb{R}^{np}$, where $\bm{\beta}(s_{i})\triangleq[\beta_{1}(s_{i}),...,\beta_{p}(s_{i})]^{\top}\in\mathbb{R}^{p}$. Then we define the matrix $\tilde{\bm{H}}$ and reparameterize coefficient matrix accordingly
$\tilde{\bm{H}}\triangleq\begin{bmatrix}\pi\bm{H}\\ \frac{\lambda_{2}}{\lambda_{1}}\bm{I}\end{bmatrix}\in\mathbb{R}^{n\times n},~~\bm{\theta}_{k}=\begin{bmatrix}\pi\bm{H}\\ \frac{\lambda_{2}}{\lambda_{1}}\bm{I}\end{bmatrix}\bm{\beta}_{k}=\tilde{\bm{H}}\bm{\beta}_{k},$ (12)
where $\bm{\theta}_{k}\triangleq[\theta_{k}(s_{1}),...,\theta_{k}(s_{n})]^{\top}\in\mathbb{R}^{n}$ is the parameter vector and $\bm{I}$ denotes the . Since $\tilde{\bm{H}}$ is invertible, we reparameterize $\bm{\beta}$ as equation (12) and express $\bm{\beta}_{k}$ as $\bm{\beta}_{k}=\tilde{\bm{H}}^{-1}\bm{\theta}_{k}$. Based on this transformation, we reparameterize the design matrix as
$\tilde{\bm{X}}\triangleq\left[\text{diag}(\bm{x}_{1})\tilde{\bm{H}}^{-1},...,\text{diag}(\bm{x}_{p})\tilde{\bm{H}}^{-1}\right].$ (13)
$L_{\text{FLAT}}(\bm{\theta})=\frac{1}{2}\|\mathbf{y}-\tilde{\bm{X}}\bm{\theta}\|_{2}^{2}+\lambda_{1}\sum_{k=1}^{p}\|\bm{\theta}_{k}\|_{1},$ (14)
where $\bm{\theta}\triangleq[\bm{\theta}_{1}^{\top},...,\bm{\theta}_{p}^{\top}]^{\top}\in\mathbb{R}^{np\times 1}$. We define the FLAT estimator of the coefficients as $\hat{\bm{\theta}}_{\mathrm{FLAT}}=\arg\min_{\bm{\theta}\in\Theta}L_{\text{FLAT}}(\bm{\theta})$. We can obtain the FLAT estimator of original regression coefficients using $\hat{\bm{\beta}}_{\mathrm{FLAT}}=\tilde{\bm{H}}^{-1}\hat{\bm{\theta}}_{\mathrm{FLAT}}$.
By minimizing (14), $\hat{\bm{\theta}}_{\mathrm{FLAT}}$ can detect the discrete spatial heterogeneity in effects while estimating the continuous spatial heterogeneity, and also select high-dimensional spatial features. As $L_{1}$ penalty terms are not differentiable at the point 0, the standard soft-thresholding can be used to derive the updates in optimization algorithm [Tibshirani, 1996, Xing et al., 2025]. To indicate non-convex spatial clusters in terms of feature effects, some classic density-based clustering algorithm [Ester et al., 1996, Miller and Han, 2009] can be used to further obtain the spatial surface in specific applications, such as the inverse temperature-salinity surface in thermohaline circulation.
Algorithm 1 FLAT Algorithm with Adaptive-minimum-spanning Tree
2: Initialization of pairwise distances via (3) and obtain the adaptive weights $\pi_{(s_{i},s_{j})}$ in (5.1)
3: Obtain the fused structure $\bm{H}$ in (4) based on edge set $\mathbb{E}$, which is built by MST
from Prim’s algorithm
6: repeat
7:    for $j=1$ to $np$ do
8:     Update $\theta_{j}^{(t+1)}=\mathcal{S}_{\lambda_{1}}\{\theta_{j}^{(t)}-\eta\sum_{i=1}^{n}\tilde{X}_{ij}(y_{i}-\hat{y}_{i}^{(t)}+\tilde{X}_{ij}\theta_{j}^{(t)})\}$
    where $\mathcal{S}_{\lambda}(x)=\text{sign}(x)(|x|-\lambda)_{+}$ is soft-thresholding operator
9:    end for
10: until $\|\bm{\theta}^{(t+1)}-\bm{\theta}^{(t)}\|_{2}<10^{-6}$
13: for each $(\epsilon,minPts)$ do
14:    Initialize all locations as unvisited
15:    for each unvisited location $s_{i}$ do
17:     if $|N_{\epsilon}(s_{i})|\geq minPts$ then
18:      Create new cluster: $cluster.id\leftarrow cluster.id+1$
19:      for each $s_{j}\in N_{\epsilon}(s_{i})$ do
20:       Compute $N_{\epsilon}(s_{j})$
21:       if $|N_{\epsilon}(s_{j})|\geq minPts$ then
22:        Merge $N_{\epsilon}(s_{j})$ into $N_{\epsilon}(s_{i})$
23:       else
24:        Mark $s_{j}$ as noise
25:       end if
26:      end for
27:      Expand cluster: Add $N_{\epsilon}(s_{i})$ to cluster $C_{cluster.id}$
28:      Mark all locations in $C_{cluster.id}$ as visited
29:     end if
30:    end for
31:    Evaluate and select the clustering labels
32: end for
5.2 The evaluations of clustering
Many evaluations are widely used to measure the results of clustering, such as the rand Index and adjusted Rand index with known clustering labels, and Silhouette coefficient and Calinski-Harabasz index without true labels. In this section, we will introduce the these measures and more discussion and usage can be find in Xing et al. [2024].